### v1

# RMT: Retentive Networks Meet Vision Transformers

Abstract
Transformer first appears in the field of natural language processing and is later migrated to the computer vision domain, where it demonstrates excellent performance in vision tasks. However, recently, Retentive Network (RetNet) has emerged as an architecture with the potential to replace Transformer, attracting widespread attention in the NLP community. Therefore, we raise the question of whether transferring RetNet’s idea to vision can also bring outstanding performance to vision tasks. To address this, we combine RetNet and Transformer to propose RMT. Inspired by RetNet, RMT introduces explicit decay into the vision backbone, bringing prior knowledge related to spatial distances to the vision model. This distance-related spatial prior allows for explicit control of the range of tokens that each token can attend to. Additionally, to reduce the computational cost of global modeling, we decompose this modeling process along the two coordinate axes of the image. Abundant experiments have demonstrated that our RMT exhibits exceptional performance across various computer vision tasks. For example, RMT achieves 84.1% Top1-acc on ImageNet-1k using merely 4.5G FLOPs. To the best of our knowledge, among all models, RMT achieves the highest Top1-acc when models are of similar size and trained with the same strategy. Moreover, RMT significantly outperforms existing vision backbones in downstream tasks such as object detection, instance segmentation, and semantic segmentation. Our work is still in progress.
1 Introduction
Model #Params Top1 Acc.
MaxViT-T [28] 31M 83.6
SMT-S [31] 20M 83.7
BiFormer-S [66] 26M 83.8
RMT-S (Ours) 27M 84.1
BiFormer-B [66] 57M 84.3
MaxViT-S [26] 69M 84.5
RMT-B (Ours) 54M 85.0
SMT-L [31] 81M 84.6
MaxViT-B [44] 120M 84.9
RMT-L (Ours) 95M 85.5
Figure 1: FLOPs vs. Top-1 accuracy on ImageNet-1K with $224\times 224$ resolution.
Since the Transformer was proposed in the field of NLP[45], it has achieved outstanding performance in many downstream tasks. Despite the modality gap between computer vision and natural language processing, researchers have successfully migrated this architecture to vision tasks, bringing us once again a huge surprise like in previous NLP tasks[10, 2, 23, 11, 28].
Recently, several powerful architectures [41, 36] have emerged in the NLP community, yet no work has attempted to transfer these NLP architectures to visual tasks. Compared to Transformer, Retentive Network (RetNet) [41] has demonstrated stronger performance on a range of NLP tasks. Therefore, we hope to be able to transfer this robust NLP architecture, RetNet, to the domain of vision.
The fundamental operator in RetNet is retention. A significant difference between retention and the basic Self-Attention operator in Transformer is the introduction of decay coefficients, which explicitly control the attention weights of each token with respect to its neighboring tokens, ensuring that the attention weights decay as the distance between tokens increases. This decay effectively introduces prior knowledge about one-dimensional distance into the model, resulting in improved performance.
Based on the findings in RetNet, we attempt to further improve retention into a 2D form and introduce it to visual tasks. Specifically, the original version of retention undergoes unidirectional decay, which suits the causal properties of NLP. However, for images without causal properties, this unidirectional decay is not suitable. Therefore, we first expand retention from unidirectional to bidirectional. Additionally, the original version of retention, designed for one-dimensional sequential information, is not appropriate for use in two-dimensional space. Hence, considering the spatial characteristics of two dimensions, we design a decay matrix based on 2D distance. Finally, to address the high computational load caused by a large number of tokens during the early stage of the vision backbone, we decompose the 2D computation process separately along the two axes of the image. We name this mechanism adapted to images as the Retentive Self-Attention (ReSA) mechanism. Based on the ReSA mechanism, we construct the RMT family.
We demonstrate the effectiveness of the proposed method through extensive experiments. As shown in Fig. 1, our RMT significantly outperforms the state-of-the-art (SOTA) models on image classification tasks. Additionally, our model exhibits more prominent advantages compared to other models in tasks such as object detection and instance segmentation. Our contributions can be summarized as follows:
  • We extend the core mechanism of the Retentive Network, retention, to the two-dimensional scenario, introducing spatial prior knowledge related to distances into vision models. The new mechanism is called Retentive Self-Attention (ReSA).
  • We decompose ReSA along two image axes, reducing computational complexity. This decomposition method effectively minimizes the computational burden while having minimal impact on the model’s performance.
  • Extensive experiments demonstrate the excellent performance of RMT. Particularly in downstream tasks such as object detection and instance segmentation, RMT exhibits significant advantages.
2 Related Work
Transformer.
Transformer architecture was firstly proposed in [45] to address the training limitation of recurrent model and then achieve massive success in many NLP tasks. By splitting the image into small, non-overlapped patches sequence, Vision Transformer (ViTs) [10] also have attracted great attention and become widely used on vision tasks [59, 17, 51, 14, 35, 4]. Unlike in the past, where RNNs and CNNs have respectively dominated the NLP and CV fields, the transformer architecture has shined through in various modalities and fields [34, 53, 37, 24].
Prior Knowledge in Transformer.
Numerous attempts have been made to incorporate prior knowledge into the Transformer model to enhance its performance. The original Transformers [10, 45] use trigonometric position encoding to provide positional information for each token. [32] proposes the use of relative positional encoding as a replacement for the original absolute positional encoding. [5] points out that zero padding in convolutional layers could also provide positional awareness for the Transformer, and this method of position encoding is highly efficient. In many studies, ConvFFN [15, 47, 13, 12] has been employed to further enrich the positional information in the Transformer. Furthermore, in the recent Retentive Network [41], explicit attenuation has been introduced to provide the model with prior knowledge based on distance changes.
Retentive Network.
Retentive Network [41] proposes the retention mechanism for sequence modeling. Compared to traditional Transformer based models [45, 10, 32, 9, 23], retention proposed in RetNet uses the explicit decay to model the prior of 1D distance. It includes three computation paradigms, i.e., parallel, recurrent, and chunkwise recurrent. In retention, it uses a decay matrix multiplied by a weight matrix to control the proportion of each token seeing its surrounding tokens based on distance priors. We attempt to extend this idea to 2D space.
3 Methodology
Figure 2: Overall architecture of RMT.
3.1 Preliminary
Retentive Network.
Retentive Network (RetNet) is a powerful architecture for language models. This work proposes the retention mechanism for sequence modeling. Retention brings the explicit decay to the language model, which Transformers do not have. Retention firstly considers a sequence modeling problem in a recurrent manner. It can be written as Eq. 1:
$o_{n}=\sum_{m=1}^{n}\gamma^{n-m}(Q_{n}e^{in\theta})(K_{m}e^{im\theta})^{\dagger}v_{m}$ (1)
During training, for a parallel training process, Eq. 1 is writed as Eq. 2:
$Q=(XW_{Q})\odot\Theta,$ $\quad K=(XW_{K})\odot\overline{\Theta},\quad V=XW_{V}$ (2)
$\Theta_{n}=e^{in\theta},$ $\quad D_{nm}=\left\{\begin{aligned} &\gamma^{n-m},&n\geq m\\ &0,&n<m\\ \end{aligned}\right.$
$\mathrm{Rete}$ $\mathrm{ntion}(X)=(QK^{\intercal}\odot D)V$
where $\overline{\Theta}$ is the complex conjugate of $\Theta$, and $D\in\mathbb{R}^{|x|\times|x|}$ contains both causal masking and exponential decay, which symbolize the relative distance in one-dimensional sequence, which brings the prior knowledge. Based on the 1D explicit decay in the retention, we try to develop it to 2D and bring the spatial prior knowledge to the vision model.
3.2 Retentive Self-Attention
Unidirectional to Bidirectional.
Due to the causal nature of language tasks, the retention in RetNet is unidirectional, meaning that each token can only attend to the tokens preceding it and not those following it. This is not suitable for tasks without causal properties, such as image recognition tasks. Therefore, we first extend the retention to two dimensions, where for each token, its output becomes Eq. 3:
$o_{n}=\sum_{m=1}^{N}\gamma^{|n-m|}(Q_{n}e^{in\theta})(K_{m}e^{im\theta})^{\dagger}v_{m}$ (3)
where $N$ is the number of tokens. The equation can be rearranged into a parallel form, expressed as Eq. 4:
$\mathrm{BiRetention}(X)$ $=(QK^{\intercal}\odot D^{Bi})V$ (4)
$D_{nm}^{Bi}$ $=\gamma^{|n-m|}$
where $\mathrm{BiRetention}$ denotes the retention with bidirectional modeling ability.
1D to 2D.
Although retention now has the ability for bi-directional modeling, this modeling capability remains limited to a one-dimensional level and is still not applicable to two-dimensional images. Therefore, we further extend the one-dimensional retention to two dimensions.
For images, each token has a unique two-dimensional coordinate within the plane. For the $n$th token, we use $(x_{n},y_{n})$ to represent its two-dimensional coordinate. Based on the 2D coordinates of each token, we modify each element in the matrix $D$ to be the Manhattan distance between the corresponding token pairs at their respective positions, completing the transformation from a 1D to a 2D decay coefficient. The matrix $D$ transfers to Eq. 5:
$D_{nm}^{2d}=\gamma^{|x_{n}-x_{m}|+|y_{n}-y_{m}|}$ (5)
In the retention [41], the $\mathrm{Softmax}$ is abandoned and replaced with a gating function to increase the nonlinearity of the operator. However, according to our experiments, this approach does not yield better results for vision models. Instead, it introduces additional parameters and computational complexity. Therefore, we still use $\mathrm{Softmax}$ to introduce nonlinearity to our model. Based on the steps mentioned above, our Retentive Self-Attention can be expressed as Eq. 6:
$\mathrm{ReSA}(X)$ $=\mathrm{Softmax}(QK^{\intercal}\odot D^{2d})V$ (6)
$D_{nm}^{2d}$ $=\gamma^{|x_{n}-x_{m}|+|y_{n}-y_{m}|}$
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
tiny model $\sim 2.5$G PVTv2-b1 [47] 13 2.1 78.7
QuadTree-B-b1 [42] 14 2.3 80.0
RegionViT-T [2] 14 2.4 80.4
MPViT-XS [26] 11 2.9 80.9
tiny-MOAT-2 [55] 10 2.3 81.0
VAN-B1 [16] 14 2.5 81.1
BiFormer-T [66] 13 2.2 81.4
CrossFormer-T [48] 28 2.9 81.5
NAT-M [18] 20 2.7 81.8
QnA-T [1] 16 2.5 82.0
SMT-T [31] 12 2.4 82.2
RMT-T 14 2.5 82.4
small model $\sim 4.5$G DeiT-S [43] 22 4.6 79.9
Swin-T [32] 29 4.5 81.3
ConvNeXt-T [33] 29 4.5 82.1
Focal-T [56] 29 4.9 82.2
InceptionNeXt-T [60] 28 4.2 82.3
FocalNet-T [57] 29 4.5 82.3
RegionViT-S [2] 31 5.3 82.6
CSWin-T [9] 23 4.3 82.7
MPViT-S [26] 23 4.7 83.0
ScalableViT-S [58] 32 4.2 83.1
MOAT-0 [55] 28 5.7 83.3
Ortho-S [22] 24 4.5 83.4
InternImage-T [49] 30 5.0 83.5
CMT-S [15] 25 4.0 83.5
FAT-B3 [13] 29 4.4 83.6
MaxViT-T [44] 31 5.6 83.6
SMT-S [31] 20 4.8 83.7
BiFormer-S [66] 26 4.5 83.8
RMT-S 27 4.5 84.1
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
base model $\sim 9.0$G Swin-S [32] 50 8.7 83.0
ConvNeXt-S [33] 50 8.7 83.1
CrossFormer-B [48] 52 9.2 83.4
InceptionNeXt-S [60] 49 8.4 83.5
PVTv2-b4 [47] 63 10.0 83.6
NAT-S [18] 51 7.8 83.7
Quadtree-B-b4 [42] 64 11.5 84.0
Ortho-B [22] 50 8.6 84.0
ScaleViT-B [58] 81 8.6 84.1
MOAT-1 [55] 42 9.1 84.2
InternImage-S [49] 50 8.0 84.2
BiFormer-B [66] 57 9.8 84.3
MViTv2-B [28] 52 10.2 84.4
CMT-B [15] 46 9.3 84.5
iFormer-B [40] 48 9.4 84.6
RMT-B 54 9.7 85.0
large model $\sim 18.0$G DeiT-B [43] 86 17.5 81.8
Swin-B [32] 88 15.4 83.3
LITv2 [35] 87 13.2 83.6
CrossFormer-L [48] 92 16.1 84.0
Ortho-L [22] 88 15.4 84.2
CSwin-B [9] 78 15.0 84.2
MPViT-B [26] 75 16.4 84.3
ScalableViT-L [58] 104 14.7 84.4
SMT-L [31] 81 17.7 84.6
MOAT-2 [55] 73 17.2 84.7
iFormer-L [40] 87 14.0 84.8
CMT-L [15] 75 19.5 84.8
InterImage-B [49] 97 16.0 84.9
MaxViT-B [44] 120 23.4 84.9
RMT-L 95 18.2 85.5
Table 1: Comparison with the state-of-the-art on ImageNet-1K classification.
Decomposed ReSA in Early Stage.
The current ReSA is not entirely applicable to image recognition tasks. This is because, in the early stages of the vision backbone, there are a large number of tokens, resulting in excessive computational costs for Attention. This is also the problem that most variants of Vision Transformers strive to solve [63, 54, 59, 50, 51, 39].
Our ReSA also encounters this problem. Therefore, we decompose ReSA into two axes of the image, as described in the specific process shown in Eq. 7:
$Q_{H},K_{H}=(Q,K)^{\mathrm{B,L,C->B,W,H,C}}$ (7)
$Q_{W},K_{W}=(Q,K)^{\mathrm{B,L,C->B,H,W,C}}$
$Attn_{H}=\mathrm{Softmax}(Q_{H}K_{H}^{\intercal}\odot D^{H})$
$Attn_{W}=\mathrm{Softmax}(Q_{W}K_{W}^{\intercal}\odot D^{W})$
$D^{H}_{nm}=\gamma^{|y_{n}-y_{m}|},D^{W}_{nm}=\gamma^{|x_{n}-x_{m}|}$
$\mathrm{ReSA_{dec}}(X)=Attn_{H}(Attn_{W}V)$
Based on this decomposition of ReSA, the shape of the receptive field of each token is shown in Fig. 3, which is identical to the shape of the complete ReSA’s receptive field.
Figure 3: Illustration of decomposed ReSA.
In order to further enhance the local expression capability of ReSA, we also introduce a local enhancement module using DWConv:
$X_{out}=\mathrm{ReSA}(X)+\mathrm{LCE}(X);$ (8)
3.3 Overall Architecture
The architecture of our entire model is shown in Fig 2. Similar to traditional backbones, it is divided into four stages. The first three stages utilize the decomposed ReSA, while the last stage uses the original ReSA. Like many previous backbones [15, 66, 27], we incorporate CPE [5] into our model.
4 Experiments
Backbone Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP^{b}_{M}$ $AP^{b}_{L}$
PVT-T [46] 33 240 39.8 62.2 43.0 37.4 59.3 39.9 23 221 39.4 59.8 42.0 25.5 42.0 52.1
PVTv2-B1 [47] 33 243 41.8 54.3 45.9 38.8 61.2 41.6 23 225 41.2 61.9 43.9 25.4 44.5 54.3
MPViT-XS [26] 30 231 44.2 66.7 48.4 40.4 63.4 43.4 20 211 43.8 65.0 47.1 28.1 47.6 56.5
FAT-B2 [13] 33 215 45.2 67.9 49.0 41.3 64.6 44.0 23 196 44.0 65.2 47.2 27.5 47.7 58.8
RMT-T 33 218 47.1 68.8 51.7 42.6 65.8 45.9 23 199 45.1 66.2 48.1 28.8 48.9 61.1
ResNet-50 [19] 44 260 38.0 58.6 41.4 34.4 55.1 36.7 38 239 36.3 55.3 38.6 19.3 40.4 48.8
Swin-T [32] 48 267 43.7 66.6 47.7 39.8 63.3 42.7 38 248 41.7 63.1 44.3 27.0 45.3 54.7
CMT-S [15] 45 249 44.6 66.8 48.9 40.7 63.9 43.4 44 231 44.3 65.5 47.5 27.1 48.3 59.1
CrossFormer-S [48] 50 301 45.4 68.0 49.7 41.4 64.8 44.6 41 272 44.4 65.8 47.4 28.2 48.4 59.4
ScalableViT-S [58] 46 256 45.8 67.6 50.0 41.7 64.7 44.8 36 238 45.2 66.5 48.4 29.2 49.1 60.3
MPViT-S [26] 43 268 46.4 68.6 51.2 42.4 65.6 45.7 32 248 45.7 57.3 48.8 28.7 49.7 59.2
CSWin-T [9] 42 279 46.7 68.6 51.3 42.2 65.6 45.4
InternImage-T [49] 49 270 47.2 69.0 52.1 42.5 66.1 45.8
SMT-S [31] 40 265 47.8 69.5 52.1 43.0 66.6 46.1
BiFormer-S [66] 47.8 69.8 52.3 43.2 66.8 46.5 45.9 66.9 49.4 30.2 49.6 61.7
RMT-S 46 262 49.0 70.8 53.9 43.9 67.8 47.4 36 244 47.8 69.1 51.8 32.1 51.8 63.5
ResNet-101 [19] 63 336 40.4 61.1 44.2 36.4 57.7 38.8 58 315 38.5 57.8 41.2 21.4 42.6 51.1
Swin-S [32] 69 359 45.7 67.9 50.4 41.1 64.9 44.2 60 339 44.5 66.1 47.4 29.8 48.5 59.1
ScalableViT-B [58] 95 349 46.8 68.7 51.5 42.5 65.8 45.9 85 330 45.8 67.3 49.2 29.9 49.5 61.0
InternImage-S [49] 69 340 47.8 69.8 52.8 43.3 67.1 46.7
CSWin-S [9] 54 342 47.9 70.1 52.6 43.2 67.1 46.2
BiFormer-B [66] 48.6 70.5 53.8 43.7 67.6 47.1 47.1 68.5 50.4 31.3 50.8 62.6
RMT-B 73 373 51.1 72.5 56.1 45.5 69.7 49.3 63 355 49.1 70.3 53.0 32.9 53.2 64.2
Swin-B [32] 107 496 46.9 69.2 51.6 42.3 66.0 45.5 98 477 45.0 66.4 48.3 28.4 49.1 60.6
PVTv2-B5 [47] 102 557 47.4 68.6 51.9 42.5 65.7 46.0
Focal-B [56] 110 533 47.8 70.2 52.5 43.2 67.3 46.5 101 514 46.3 68.0 49.8 31.7 50.4 60.8
MPViT-B [26] 95 503 48.2 70.0 52.9 43.5 67.1 46.8 85 482 47.0 68.4 50.8 29.4 51.3 61.5
CSwin-B [9] 97 526 48.7 70.4 53.9 43.9 67.8 47.3
InternImage-B [49] 115 501 48.8 70.9 54.0 44.0 67.8 47.4
RMT-L 114 557 51.6 73.1 56.5 45.9 70.3 49.8 104 537 49.4 70.6 53.1 34.2 53.9 65.2
Table 2: Comparison to other backbones using RetinaNet and Mask R-CNN on COCO val2017 object detection and instance segmentation.
Backbone Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
Swin-T [32] 48 267 46.0 68.1 50.3 41.6 65.1 44.9
SMT-S [31] 40 265 49.0 70.1 53.4 43.4 67.3 46.7
CSWin-T [9] 42 279 49.0 70.7 53.7 43.6 67.9 46.6
InternImage-T [49] 49 270 49.1 70.4 54.1 43.7 67.3 47.3
RMT-S 46 262 50.7 71.9 55.6 44.9 69.1 48.4
ConvNeXt-S [33] 70 348 47.9 70.0 52.7 42.9 66.9 46.2
InternImage-S [49] 69 340 49.7 71.1 54.5 44.5 68.5 47.8
SMT-B [31] 52 328 49.8 71.0 54.4 44.0 68.0 47.3
CSWin-S [9] 54 342 50.0 71.3 54.7 44.5 68.4 47.7
RMT-B 73 373 52.2 72.9 57.0 46.1 70.4 49.9
Table 3: Comparison to other backbones using Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
We conducted extensive experiments on multiple vision tasks, such as image classification on ImageNet-1K [8], object detection and instance segmentation on COCO 2017 [30], and semantic segmentation on ADE20K [65]. We also make ablation studies to validate the importance of each component in RMT.
4.1 Image Classification
Settings.
We train our models on ImageNet-1K [8] from scratch. And we follow the same training strategy in DeiT [43] for a fair comparison. The maximum rates of increasing stochastic depth [21] are set to 0.1/0.15/0.4/0.5 for RMT-T/S/B/L [21], respectively. We use the AdamW optimizer with a cosine decay learning rate scheduler to train the models. We set the initial learning rate, weight decay, and batch size to 0.001, 0.05, and 1024, respectively. We adopt the strong data augmentation and regularization used in [32]. Our settings are RandAugment [7] (randm9-mstd0.5-inc1), Mixup [62] (prob=0.8), CutMix [61] (prob=1.0), Random Erasing [64] (prob=0.25).
Results.
We compare RMT against many state-of-the-art models in Tab. 1. Results in the table demonstrate that RMT consistently outperforms previous models across all settings. Specifically, RMT-S achieves 84.1% Top1-accuracy with only 4.5 GFLOPs. RMT-B also surpasses iFormer [40] by 0.4% with similar FLOPs. Furthermore, our RMT-L model surpasses MaxViT-B [44] in top1-accuracy by 0.6% while using fewer FLOPs. Our RMT-T has also outperformed many lightweight models.
4.2 Object Detection and Instance Segmentation
Settings.
We adopt MMDetection [3] to implement RetinaNet [29] and Mask-RCNN [20]. We use the commonly used ”1×1$\times$1 ×” (12 training epochs) setting for the two strategies and ”$3\times+\mathrm{MS}$” for Mask-RCNN. Following [32], during training, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. We adopt the AdamW optimizer with a learning rate of 0.0001, weight decay of 0.05, and batch size of 16 to optimize the model. For ”1×1$\times$1 ×” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 8 and 11. While for ”$3\times+\mathrm{MS}$” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 27 and 33.
Results.
Tab. 2 and Tab. 3 shows the results with RetinaNet and Mask R-CNN. The results demonstrate that our RMT performs best in all comparisons. For the RetinaNet framework, our RMT-T outperforms FAT-B2 by +1.1 AP, while S/B/L also perform better than other methods. As for the Mask R-CNN with ”1×1$\times$1 ×” schedule, RMT-L outperforms the recent InternImage-B by +1.8 box AP and +1.9 mask AP. For ”$3\times+\mathrm{MS}$” schedule, RMT-S outperforms InternImage-T for +1.6 box AP and +1.2 mask AP. All the above results tell that RMT outperforms its counterparts by evident margins.
4.3 Semantic Segmentation
Backbone Semantic FPN
Params(M) FLOPs(G) mIoU(%)
DAT-T [51] 32 198 42.6
CrossFormer-S [48] 34 221 46.0
UniFormer-S [27] 25 247 46.6
CSWin-T [9] 26 202 48.2
Shuted-S [39] 26 183 48.2
RMT-S 30 180 49.4
DAT-S 53 320 46.1
UniFormer-B [27] 54 350 47.7
CrossFormer-B [48] 56 331 47.7
CSWin-S [9] 39 271 49.2
RMT-B 57 294 50.4
DAT-B 92 481 47.0
CrossFormer-L 95 497 48.7
CSWin-B [9] 81 464 49.9
RMT-L 98 482 51.4
Table 4: Comparison with the state-of-the-art on ADE20K. The framework is SemanticFPN.
Backbone UperNet
Params(M) FLOPs(G) mIoU(%)
DAT-T [51] 60 957 45.5
NAT-T [18] 58 934 47.1
InternImage-T [49] 59 944 47.9
HorNet-T [38] 55 924 49.2
SMT-S [31] 50 935 49.2
RMT-S 56 937 49.8
DAT-S [51] 81 1079 48.3
SMT-B [31] 62 1004 49.6
HorNet-S [38] 85 1027 50.0
InterImage-S [49] 80 1017 50.2
CSWin-S [9] 65 1027 50.4
RMT-B 83 1051 51.5
Table 5: Comparison with the state-of-the-art on ADE20K. The framework is UperNet.
Settings.
We adopt the Semantic FPN [25] and UperNet [52] based on MMSegmentation [6], apply RMTs which are pretrained on ImageNet-1K as backbone. We use the same setting of PVT [46] to train the Semantic FPN, and we train the model for 80k iterations. All models are trained with the input resolution of $512\times 512$. When testing the model, we resize the shorter side of the image to 512 pixels. As for UperNet, we follow the default settings in Swin [32]. We take AdamW with a weight decay of 0.01 as the optimizer to train the models for 160K iterations. The learning rate is set to $6\times 10^{-5}$ with 1500 iterations warmup.
Results.
The results of semantic segmentation can be found in Tab. 4 and Tab. 5. All the FLOPs are measured with the resolution of $512\times 2048$. All our models achieve the best performance in all comparisons. Specifically, our RMT-S exceeds Shunted-S for +1.2 mIoU with Semantic FPN. Moreover, our RMT-B outperforms the recent InternImage-S for +1.3 mIoU. All the above results demonstrate our model’s superiority in dense prediction.
4.4 Ablation Study
Model Params(M) FLOPs(G) Top1-acc(%)
RMT-T 14 2.5 82.4
w/o $\gamma$ 14 2.5 81.6
Table 6: Ablation study
$\gamma$ decay.
We verify the impact of explicit decay on the model, as shown in the Tab. 6. Explicit decay improves the performance of the model.
5 Conclusion
We propose RMT, a vision backbone that integrates retentive network and Vision Transformer. RMT introduces explicit decay related to distance, which brings spatial prior knowledge to visual models. The new mechanism is called Retentive Self-Attention (ReSA). To reduce the complexity of the model, RMT also employs a method that decomposes the ReSA into two axes. Extensive experiments validate the effectiveness of RMT, especially in downstream tasks such as object detection, where RMT demonstrates significant advantages. Our work is still in progress.

### v6

@@
 Abstract
-Transformer first appears in the field of natural language processing and is later migrated to the computer vision domain, where it demonstrates excellent performance in vision tasks.
-However, recently, Retentive Network (RetNet) has emerged as an architecture with the potential to replace Transformer, attracting widespread attention in the NLP community.
-Therefore, we raise the question of whether transferring RetNet’s idea to vision can also bring outstanding performance to vision tasks.
-To address this, we combine RetNet and Transformer to propose RMT.
-Inspired by RetNet, RMT introduces explicit decay into the vision backbone, bringing prior knowledge related to spatial distances to the vision model.
-This distance-related spatial prior allows for explicit control of the range of tokens that each token can attend to.
-Additionally, to reduce the computational cost of global modeling, we decompose this modeling process along the two coordinate axes of the image.
-Abundant experiments have demonstrated that our RMT exhibits exceptional performance across various computer vision tasks.
-For example, RMT achieves 84.1% Top1-acc on ImageNet-1k using merely 4.5G FLOPs.
-To the best of our knowledge, among all models, RMT achieves the highest Top1-acc when models are of similar size and trained with the same strategy.
-Moreover, RMT significantly outperforms existing vision backbones in downstream tasks such as object detection, instance segmentation, and semantic segmentation.
-Our work is still in progress.
+Vision Transformer (ViT) has gained increasing attention in the computer vision community in recent years.
+However, the core component of ViT, Self-Attention, lacks explicit spatial priors and bears a quadratic computational complexity, thereby constraining the applicability of ViT.
+To alleviate these issues, we draw inspiration from the recent Retentive Network (RetNet) in the field of NLP, and propose RMT, a strong vision backbone with explicit spatial prior for general purposes.
+Specifically, we extend the RetNet’s temporal decay mechanism to the spatial domain, and propose a spatial decay matrix based on the Manhattan distance to introduce the explicit spatial prior to Self-Attention.
+Additionally, an attention decomposition form that adeptly adapts to explicit spatial prior is proposed, aiming to reduce the computational burden of modeling global information without disrupting the spatial decay matrix.
+Based on the spatial decay matrix and the attention decomposition form, we can flexibly integrate explicit spatial prior into the vision backbone with linear complexity.
+Extensive experiments demonstrate that RMT exhibits exceptional performance across various vision tasks.
+Specifically, without extra training data, RMT achieves 84.8% and 86.1% top-1 acc on ImageNet-1k with 27M/4.5GFLOPs and 96M/18.2GFLOPs.
+For downstream tasks, RMT achieves 54.5 box AP and 47.2 mask AP on the COCO detection task, and 52.8 mIoU on the ADE20K semantic segmentation task.
+Code is available at https://github.com/qhfan/RMT
 1 Introduction
@@
 RMT-S (Ours) 27M 84.1
+RMT-S* (Ours) 27M 84.8
 BiFormer-B [75] 57M 84.3
@@
 RMT-B (Ours) 54M 85.0
+RMT-B* (Ours) 55M 85.6
 SMT-L [34] 81M 84.6
@@
 RMT-L (Ours) 95M 85.5
-Figure 1: FLOPs vs. Top-1 accuracy on ImageNet-1K with $224\times 224$ resolution.
-Since the Transformer was proposed in the field of NLP[45], it has achieved outstanding performance in many downstream tasks.
-Despite the modality gap between computer vision and natural language processing, researchers have successfully migrated this architecture to vision tasks, bringing us once again a huge surprise like in previous NLP tasks[10, 2, 23, 11, 28].
-Recently, several powerful architectures [41, 36] have emerged in the NLP community, yet no work has attempted to transfer these NLP architectures to visual tasks.
-Compared to Transformer, Retentive Network (RetNet) [41] has demonstrated stronger performance on a range of NLP tasks.
-Therefore, we hope to be able to transfer this robust NLP architecture, RetNet, to the domain of vision.
-The fundamental operator in RetNet is retention.
-A significant difference between retention and the basic Self-Attention operator in Transformer is the introduction of decay coefficients, which explicitly control the attention weights of each token with respect to its neighboring tokens, ensuring that the attention weights decay as the distance between tokens increases.
-This decay effectively introduces prior knowledge about one-dimensional distance into the model, resulting in improved performance.
-Based on the findings in RetNet, we attempt to further improve retention into a 2D form and introduce it to visual tasks.
-Specifically, the original version of retention undergoes unidirectional decay, which suits the causal properties of NLP.
-However, for images without causal properties, this unidirectional decay is not suitable.
-Therefore, we first expand retention from unidirectional to bidirectional.
-Additionally, the original version of retention, designed for one-dimensional sequential information, is not appropriate for use in two-dimensional space.
-Hence, considering the spatial characteristics of two dimensions, we design a decay matrix based on 2D distance.
-Finally, to address the high computational load caused by a large number of tokens during the early stage of the vision backbone, we decompose the 2D computation process separately along the two axes of the image.
-We name this mechanism adapted to images as the Retentive Self-Attention (ReSA) mechanism.
-Based on the ReSA mechanism, we construct the RMT family.
+RMT-L* (Ours) 96M 86.1
+Figure 1: FLOPs v.s. Top-1 accuracy on ImageNet-1K with $224\times 224$ resolution. “*” indicates the model trained with token labeling [27].
+Vision Transformer (ViT) [12] is an excellent visual architecture highly favored by researchers.
+However, as the core module of ViT, Self-Attention’s inherent structure lacking explicit spatial priors.
+Besides, the quadratic complexity of Self-Attention leads to significant computational costs when modeling global information.
+These issues limit the application of ViT.
+Many works have previously attempted to alleviate these issues [30, 35, 50, 13, 57, 16, 61].
+For example, in Swin Transformer [35], the authors partition the tokens used for self-attention by applying windowing operations.
+This operation not only reduces the computational cost of self-attention but also introduces spatial priors to the model through the use of windows and relative position encoding.
+In addition to it, NAT [19] changes the receptive field of Self-Attention to match the shape of convolution, reducing computational costs while also enabling the model to perceive spatial priors through the shape of its receptive field.
+Figure 2: Comparison among different Self-Attention mechanisms.
+In MaSA, darker colors represent smaller spatial decay rates, while lighter colors represent larger ones.
+The spatial decay rates that change with distance provide the model with rich spatial priors.
+Different from previous methods, we draw inspiration from the recently successful Retentive Network (RetNet) [46] in the field of NLP.
+RetNet utilizes a distance-dependent temporal decay matrix to provide explicit temporal prior for one-dimensional and unidirectional text data.
+ALiBi [41], prior to RetNet, also applied a similar approach and succeeded in NLP tasks.
+We extend this temporal decay matrix to the spatial domain, developing a two-dimensional bidirectional spatial decay matrix based on the Manhattan distance among tokens.
+In our space decay matrix, for a target token, the farther the surrounding tokens are, the greater the degree of decay in their attention scores.
+This property allows the target token to perceive global information while simultaneously assigning different levels of attention to tokens at varying distances.
+We introduce explicit spatial prior to the vision backbone using this spatial decay matrix.
+We name this Self-Attention mechanism, which is inspired by RetNet and incorporates the Manhattan distance as the explicit spatial prior, as Manhattan Self-Attention (MaSA).
+Besides explicit spatial priors, another issue caused by global modeling with Self-Attention is the enormous computational burden.
+Previous sparse attention mechanisms [11, 35, 53, 63, 75] and the way retention is decomposed in RetNet [46] mostly disrupt the spatial decay matrix, making them unsuitable for MaSA.
+In order to sparsely model global information without compromising the spatial decay matrix, we propose a method to decompose Self-Attention along both axes of the image.
+This decomposition method decomposes Self-Attention and the spatial decay matrix without any loss of prior information.
+The decomposed MaSA models global information with linear complexity and has the same receptive field shape as the original MaSA.
+We compare MaSA with other Self-Attention mechanisms in Fig.
+2.
+It can be seen that our MaSA introduces richer spatial priors to the model than its counterparts.
+Based on MaSA, we construct a powerful vision backbone called RMT.
 We demonstrate the effectiveness of the proposed method through extensive experiments.
 As shown in Fig.
-1, our RMT significantly outperforms the state-of-the-art (SOTA) models on image classification tasks.
-Additionally, our model exhibits more prominent advantages compared to other models in tasks such as object detection and instance segmentation.
+1, our RMT outperforms the state-of-the-art (SOTA) models on image classification tasks.
+Additionally, our model exhibits more prominent advantages compared to other models in tasks such as object detection, instance segmentation, and semantic segmentation.
 Our contributions can be summarized as follows:
-  • We extend the core mechanism of the Retentive Network, retention, to the two-dimensional scenario, introducing spatial prior knowledge related to distances into vision models.
-The new mechanism is called Retentive Self-Attention (ReSA).
-  • We decompose ReSA along two image axes, reducing computational complexity.
-This decomposition method effectively minimizes the computational burden while having minimal impact on the model’s performance.
-  • Extensive experiments demonstrate the excellent performance of RMT. Particularly in downstream tasks such as object detection and instance segmentation, RMT exhibits significant advantages.
+  • We propose a spatial decay matrix based on Manhattan distance to augment Self-Attention, creating the Manhattan Self-Attention (MaSA) with an explicit spatial prior.
+  • We propose a decomposition form for MaSA, enabling linear complexity for global information modeling without disrupting the spatial decay matrix.
+  • Leveraging MaSA, we construct RMT, a powerful vision backbone for general purposes.
+RMT attains high top-1 accuracy on ImageNet-1k in image classification without extra training data, and excels in tasks like object detection, instance segmentation, and semantic segmentation.
 2 Related Work
@@
 Unlike in the past, where RNNs and CNNs have respectively dominated the NLP and CV fields, the transformer architecture has shined through in various modalities and fields [37, 60, 42, 26].
+In the computer vision community, many studies are attempting to introduce spatial priors into ViT to reduce the data requirements for training [6, 49, 19].
+At the same time, various sparse attention mechanisms have been proposed to reduce the computational cost of Self-Attention [53, 54, 13, 57].
 Prior Knowledge in Transformer.
@@
 The original Transformers [12, 52] use trigonometric position encoding to provide positional information for each token.
-[32] proposes the use of relative positional encoding as a replacement for the original absolute positional encoding.
-[5] points out that zero padding in convolutional layers could also provide positional awareness for the Transformer, and this method of position encoding is highly efficient.
-In many studies, ConvFFN [15, 47, 13, 12] has been employed to further enrich the positional information in the Transformer.
-Furthermore, in the recent Retentive Network [41], explicit attenuation has been introduced to provide the model with prior knowledge based on distance changes.
-Retentive Network.
-Retentive Network [41] proposes the retention mechanism for sequence modeling.
-Compared to traditional Transformer based models [45, 10, 32, 9, 23], retention proposed in RetNet uses the explicit decay to model the prior of 1D distance.
-It includes three computation paradigms, i.e., parallel, recurrent, and chunkwise recurrent.
-In retention, it uses a decay matrix multiplied by a weight matrix to control the proportion of each token seeing its surrounding tokens based on distance priors.
-We attempt to extend this idea to 2D space.
+In vision tasks, [35] proposes the use of relative positional encoding as a replacement for the original absolute positional encoding.
+[6] points out that zero padding in convolutional layers could also provide positional awareness for the ViT, and this position encoding method is highly efficient.
+In many studies, Convolution in FFN [16, 54, 13] has been employed for vision models to further enrich the positional information in the ViT.
+For NLP tasks, in the recent Retentive Network [46], the temporal decay matrix has been introduced to provide the model with prior knowledge based on distance changes.
+Before RetNet, ALiBi [41] also uses a similar temporal decay matrix.
 3 Methodology
@@
 3.1 Preliminary
-Retentive Network.
+Temporal decay in RetNet.
 Retentive Network (RetNet) is a powerful architecture for language models.
 This work proposes the retention mechanism for sequence modeling.
-Retention brings the explicit decay to the language model, which Transformers do not have.
+Retention brings the temporal decay to the language model, which Transformers do not have.
 Retention firstly considers a sequence modeling problem in a recurrent manner.
@@
 $o_{n}=\sum_{m=1}^{n}\gamma^{n-m}(Q_{n}e^{in\theta})(K_{m}e^{im\theta})^{\dagger}v_{m}$ (1)
-During training, for a parallel training process, Eq. 1 is writed as Eq. 2:
+For a parallel training process, Eq. 1 is expressed as:
 $Q=(XW_{Q})\odot\Theta,$ $\quad K=(XW_{K})\odot\overline{\Theta},\quad V=XW_{V}$ (2)
@@
 $\mathrm{Rete}$ $\mathrm{ntion}(X)=(QK^{\intercal}\odot D)V$
-where $\overline{\Theta}$ is the complex conjugate of $\Theta$, and $D\in\mathbb{R}^{|x|\times|x|}$ contains both causal masking and exponential decay, which symbolize the relative distance in one-dimensional sequence, which brings the prior knowledge.
-Based on the 1D explicit decay in the retention, we try to develop it to 2D and bring the spatial prior knowledge to the vision model.
-3.2 Retentive Self-Attention
-Unidirectional to Bidirectional.
-Due to the causal nature of language tasks, the retention in RetNet is unidirectional, meaning that each token can only attend to the tokens preceding it and not those following it.
-This is not suitable for tasks without causal properties, such as image recognition tasks.
-Therefore, we first extend the retention to two dimensions, where for each token, its output becomes Eq.
+where $\overline{\Theta}$ is the complex conjugate of $\Theta$, and $D\in\mathbb{R}^{|x|\times|x|}$ contains both causal masking and exponential decay, which symbolizes the relative distance in one-dimensional sequence and brings the explicit temporal prior to text data.
+3.2 Manhattan Self-Attention
+Starting from the retention in RetNet, we evolve it into Manhattan Self-Attention (MaSA).
+Within MaSA, we transform the unidirectional and one-dimensional temporal decay observed in retention into bidirectional and two-dimensional spatial decay.
+This spatial decay introduces an explicit spatial prior linked to Manhattan distance into the vision backbone.
+Additionally, we devise a straightforward approach to concurrently decompose the Self-Attention and spatial decay matrix along the two axes of the image.
+From Unidirectional to Bidirectional Decay:
+In RetNet, retention is unidirectional due to the causal nature of text data, allowing each token to attend only to preceding tokens and not those following it.
+This characteristic is ill-suited for tasks lacking causal properties, such as image recognition.
+Hence, we initially broaden the retention to a bidirectional form, expressed as Eq.
 3:
-$o_{n}=\sum_{m=1}^{N}\gamma^{|n-m|}(Q_{n}e^{in\theta})(K_{m}e^{im\theta})^{\dagger}v_{m}$ (3)
-where $N$ is the number of tokens. The equation can be rearranged into a parallel form, expressed as Eq. 4:
 $\mathrm{BiRetention}(X)$ $=(QK^{\intercal}\odot D^{Bi})V$ (3)
 $D_{nm}^{Bi}$ $=\gamma^{|n-m|}$
-where $\mathrm{BiRetention}$ denotes the retention with bidirectional modeling ability.
-1D to 2D.
-Although retention now has the ability for bi-directional modeling, this modeling capability remains limited to a one-dimensional level and is still not applicable to two-dimensional images.
-Therefore, we further extend the one-dimensional retention to two dimensions.
-For images, each token has a unique two-dimensional coordinate within the plane.
-For the $n$th token, we use $(x_{n},y_{n})$ to represent its two-dimensional coordinate.
-Based on the 2D coordinates of each token, we modify each element in the matrix $D$ to be the Manhattan distance between the corresponding token pairs at their respective positions, completing the transformation from a 1D to a 2D decay coefficient.
-The matrix $D$ transfers to Eq.
-5:
+where $\mathrm{BiRetention}$ signifies bidirectional modeling.
+From One-dimensional to Two-dimensional Decay:
+While retention now supports bi-directional modeling, this capability remains confined to a one-dimensional level and is inadequate for two-dimensional images.
+To address this limitation, we extend the one-dimensional retention to encompass two dimensions.
+In the context of images, each token is uniquely positioned with a two-dimensional coordinate within the plane, denoted as $(x_{n},y_{n})$ for the $n$-th token.
+To adapt to this, we adjust each element in the matrix $D$ to represent the Manhattan distance between the respective token pairs based on their 2D coordinates.
+The matrix $D$ is redefined as follows:
 $D_{nm}^{2d}=\gamma^{|x_{n}-x_{m}|+|y_{n}-y_{m}|}$ (4)
-In the retention [41], the $\mathrm{Softmax}$ is abandoned and replaced with a gating function to increase the nonlinearity of the operator.
-However, according to our experiments, this approach does not yield better results for vision models.
-Instead, it introduces additional parameters and computational complexity.
-Therefore, we still use $\mathrm{Softmax}$ to introduce nonlinearity to our model.
-Based on the steps mentioned above, our Retentive Self-Attention can be expressed as Eq.
-6:
-$\mathrm{ReSA}(X)$ $=\mathrm{Softmax}(QK^{\intercal}\odot D^{2d})V$ (6)
+Figure 4: Spatial decay matrix in the decomposed MaSA.
+In the retention, the $\mathrm{Softmax}$ is abandoned and replaced with a gating function.
+This variation gives RetNet multiple flexible computation forms, enabling it to adapt to parallel training and recurrent inference processes.
+Despite this flexibility, when exclusively utilizing RetNet’s parallel computation form in our experiments, the necessity of retaining the gating function becomes debatable.
+Our findings indicate that this modification does not improve results for vision models; instead, it introduces extra parameters and computational complexity.
+Consequently, we continue to employ $\mathrm{Softmax}$ to introduce nonlinearity to our model.
+Combining the aforementioned steps, our Manhattan Self-Attention is expressed as
+$\mathrm{MaSA}(X)$ $=(\mathrm{Softmax}(QK^{\intercal})\odot D^{2d})V$ (5)
 $D_{nm}^{2d}$ $=\gamma^{|x_{n}-x_{m}|+|y_{n}-y_{m}|}$
+Decomposed Manhattan Self-Attention.
+In the early stages of the vision backbone, an abundance of tokens leads to high computational costs for Self-Attention when attempting to model global information.
+Our MaSA encounters this challenge as well.
+Utilizing existing sparse attention mechanisms [35, 11, 19, 53, 63], or the original RetNet’s recurrent/chunk-wise recurrent form directly, disrupts the spatial decay matrix based on Manhattan distance, resulting in the loss of explicit spatial prior.
+To address this, we introduce a simple decomposition method that not only decomposes Self-Attention but also decomposes the spatial decay matrix.
+The decomposed MaSA is represented in Eq.
+6.
+Specifically, we calculate attention scores separately for the horizontal and vertical directions in the image.
+Subsequently, we apply the one-dimensional bidirectional decay matrix to these attention weights.
+The one-dimensional decay matrix signifies the horizontal and vertical distances between tokens ($D^{H}_{nm}=\gamma^{|y_{n}-y_{m}|}$, $D^{W}_{nm}=\gamma^{|x_{n}-x_{m}|}$):
+$Attn_{H}=\mathrm{Softmax}(Q_{H}K_{H}^{\intercal})\odot D^{H},$ (6)
+$Attn_{W}=\mathrm{Softmax}(Q_{W}K_{W}^{\intercal})\odot D^{W},$
+$\mathrm{MaSA}(X)=Attn_{H}(Attn_{W}V)^{\intercal}$
+Based on the decomposition of MaSA, the shape of the receptive field of each token is shown in Fig.
+4, which is identical to the shape of the complete MaSA’s receptive field.
+Fig.
+4 indicates that our decomposition method fully preserves the explicit spatial prior.
 Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
@@
 BiFormer-T [75] 13 2.2 81.4
+Conv2Former-N [23] 15 2.2 81.5
 CrossFormer-T [55] 28 2.9 81.5
@@
 QnA-T [1] 16 2.5 82.0
+GC-ViT-XT [20] 20 2.6 82.0
 SMT-T [34] 12 2.4 82.2
@@
 Focal-T [63] 29 4.9 82.2
-InceptionNeXt-T [60] 28 4.2 82.3
 FocalNet-T [64] 29 4.5 82.3
@@
 ScalableViT-S [65] 32 4.2 83.1
+SG-Former-S [15] 23 4.8 83.2
 MOAT-0 [62] 28 5.7 83.3
@@
 CMT-S [16] 25 4.0 83.5
-FAT-B3 [13] 29 4.4 83.6
 MaxViT-T [51] 31 5.6 83.6
@@
 RMT-S 27 4.5 84.1
+LV-ViT-S* [27] 26 6.6 83.3
+UniFormer-S* [30] 24 4.2 83.4
+WaveViT-S* [66] 23 4.7 83.9
+Dual-ViT-S* [67] 25 5.4 84.1
+VOLO-D1* [68] 27 6.8 84.2
+BiFormer-S* [75] 26 4.5 84.3
+RMT-S* 27 4.5 84.8
 Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
@@
 CrossFormer-B [55] 52 9.2 83.4
-InceptionNeXt-S [60] 49 8.4 83.5
-PVTv2-b4 [47] 63 10.0 83.6
 NAT-S [19] 51 7.8 83.7
@@
 InternImage-S [56] 50 8.0 84.2
+DaViT-S [10] 50 8.8 84.2
+GC-ViT-S [20] 51 8.5 84.3
 BiFormer-B [75] 57 9.8 84.3
 MViTv2-B [31] 52 10.2 84.4
-CMT-B [15] 46 9.3 84.5
 iFormer-B [45] 48 9.4 84.6
 RMT-B 54 9.7 85.0
-large model $\sim 18.0$G DeiT-B [43] 86 17.5 81.8
-Swin-B [32] 88 15.4 83.3
+WaveViT-B* [66] 34 7.2 84.8
+UniFormer-B* [30] 50 8.3 85.1
+Dual-ViT-B* [67] 43 9.3 85.2
+BiFormer-B* [75] 58 9.8 85.4
+RMT-B* 55 9.7 85.6
+large model $\sim 18.0$G Swin-B [35] 88 15.4 83.3
+CaiT-M24 [50] 186 36 83.4
 LITv2 [39] 87 13.2 83.6
@@
 CSwin-B [11] 78 15.0 84.2
-MPViT-B [26] 75 16.4 84.3
-ScalableViT-L [58] 104 14.7 84.4
 SMT-L [34] 81 17.7 84.6
 MOAT-2 [62] 73 17.2 84.7
+SG-Former-B [15] 78 15.6 84.7
 iFormer-L [45] 87 14.0 84.8
-CMT-L [15] 75 19.5 84.8
 InterImage-B [56] 97 16.0 84.9
 MaxViT-B [51] 120 23.4 84.9
+GC-ViT-B [20] 90 14.8 85.0
 RMT-L 95 18.2 85.5
-Table 1: Comparison with the state-of-the-art on ImageNet-1K classification.
-Decomposed ReSA in Early Stage.
-The current ReSA is not entirely applicable to image recognition tasks.
-This is because, in the early stages of the vision backbone, there are a large number of tokens, resulting in excessive computational costs for Attention.
-This is also the problem that most variants of Vision Transformers strive to solve [63, 54, 59, 50, 51, 39].
-Our ReSA also encounters this problem. Therefore, we decompose ReSA into two axes of the image, as described in the specific process shown in Eq. 7:
-$Q_{H},K_{H}=(Q,K)^{\mathrm{B,L,C->B,W,H,C}}$ (7)
-$Q_{W},K_{W}=(Q,K)^{\mathrm{B,L,C->B,H,W,C}}$
-$Attn_{H}=\mathrm{Softmax}(Q_{H}K_{H}^{\intercal}\odot D^{H})$
-$Attn_{W}=\mathrm{Softmax}(Q_{W}K_{W}^{\intercal}\odot D^{W})$
-$D^{H}_{nm}=\gamma^{|y_{n}-y_{m}|},D^{W}_{nm}=\gamma^{|x_{n}-x_{m}|}$
-$\mathrm{ReSA_{dec}}(X)=Attn_{H}(Attn_{W}V)$
-Based on this decomposition of ReSA, the shape of the receptive field of each token is shown in Fig. 3, which is identical to the shape of the complete ReSA’s receptive field.
-Figure 3: Illustration of decomposed ReSA.
-In order to further enhance the local expression capability of ReSA, we also introduce a local enhancement module using DWConv:
-$X_{out}=\mathrm{ReSA}(X)+\mathrm{LCE}(X);$ (8)
+VOLO-D3* [68] 86 20.6 85.4
+WaveViT-L* [66] 58 14.8 85.5
+UniFormer-L* [30] 100 12.6 85.6
+Dual-ViT-L* [67] 73 18.0 85.7
+RMT-L* 96 18.2 86.1
+Table 1: Comparison with the state-of-the-art on ImageNet-1K classification. “*” indicates the model trained with token labeling [27].
+To further enhance the local expression capability of MaSA, following [75], we introduce a Local Context Enhancement module using DWConv:
+$X_{out}=\mathrm{MaSA}(X)+\mathrm{LCE}(V);$ (7)
 3.3 Overall Architecture
-The architecture of our entire model is shown in Fig 2.
-Similar to traditional backbones, it is divided into four stages.
-The first three stages utilize the decomposed ReSA, while the last stage uses the original ReSA.
-Like many previous backbones [15, 66, 27], we incorporate CPE [5] into our model.
-4 Experiments
+We construct the RMT based on MaSA, and its architecture is illustrated in Fig.
+3.
+Similar to previous general vision backbones [53, 54, 35, 71], RMT is divided into four stages.
+The first three stages utilize the decomposed MaSA, while the last uses the original MaSA.
+Like many previous backbones [16, 75, 30, 72], we incorporate CPE [6] into our model.
 Backbone Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
@@
 MPViT-XS [29] 30 231 44.2 66.7 48.4 40.4 63.4 43.4 20 211 43.8 65.0 47.1 28.1 47.6 56.5
-FAT-B2 [13] 33 215 45.2 67.9 49.0 41.3 64.6 44.0 23 196 44.0 65.2 47.2 27.5 47.7 58.8
 RMT-T 33 218 47.1 68.8 51.7 42.6 65.8 45.9 23 199 45.1 66.2 48.1 28.8 48.9 61.1
-ResNet-50 [19] 44 260 38.0 58.6 41.4 34.4 55.1 36.7 38 239 36.3 55.3 38.6 19.3 40.4 48.8
 Swin-T [35] 48 267 43.7 66.6 47.7 39.8 63.3 42.7 38 248 41.7 63.1 44.3 27.0 45.3 54.7
@@
 Table 2: Comparison to other backbones using RetinaNet and Mask R-CNN on COCO val2017 object detection and instance segmentation.
-Backbone Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
-$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
-Swin-T [32] 48 267 46.0 68.1 50.3 41.6 65.1 44.9
-SMT-S [31] 40 265 49.0 70.1 53.4 43.4 67.3 46.7
-CSWin-T [9] 42 279 49.0 70.7 53.7 43.6 67.9 46.6
-InternImage-T [49] 49 270 49.1 70.4 54.1 43.7 67.3 47.3
-RMT-S 46 262 50.7 71.9 55.6 44.9 69.1 48.4
-ConvNeXt-S [33] 70 348 47.9 70.0 52.7 42.9 66.9 46.2
-InternImage-S [49] 69 340 49.7 71.1 54.5 44.5 68.5 47.8
-SMT-B [31] 52 328 49.8 71.0 54.4 44.0 68.0 47.3
-CSWin-S [9] 54 342 50.0 71.3 54.7 44.5 68.4 47.7
-RMT-B 73 373 52.2 72.9 57.0 46.1 70.4 49.9
-Table 3: Comparison to other backbones using Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
+4 Experiments
 We conducted extensive experiments on multiple vision tasks, such as image classification on ImageNet-1K [9], object detection and instance segmentation on COCO 2017 [33], and semantic segmentation on ADE20K [74].
 We also make ablation studies to validate the importance of each component in RMT.
+More details can be found in Appendix.
 4.1 Image Classification
@@
 We train our models on ImageNet-1K [9] from scratch.
-And we follow the same training strategy in DeiT [43] for a fair comparison.
+We follow the same training strategy in [49], with the only supervision being classification loss for a fair comparison.
 The maximum rates of increasing stochastic depth [24] are set to 0.1/0.15/0.4/0.5 for RMT-T/S/B/L [24], respectively.
@@
 Our settings are RandAugment [8] (randm9-mstd0.5-inc1), Mixup [70] (prob=0.8), CutMix [69] (prob=1.0), Random Erasing [73] (prob=0.25).
+In addition to the conventional training methods, similar to LV-ViT [27] and VOLO [68], we train a model that utilizes token labeling to provide supplementary supervision.
 Results.
@@
 Our RMT-T has also outperformed many lightweight models.
+As for the model trained using token labeling, our RMT-S outperforms the current state-of-the-art BiFormer-S by 0.5%.
 4.2 Object Detection and Instance Segmentation
+Backbone Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
+$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
+ConvNeXt-T [36] 48 262 46.2 67.9 50.8 41.7 65.0 45.0
+Focal-T [63] 49 291 47.2 69.4 51.9 42.7 66.5 45.9
+NAT-T [19] 48 258 47.8 69.0 52.6 42.6 66.0 45.9
+GC-ViT-T [20] 48 291 47.9 70.1 52.8 43.2 67.0 46.7
+MPViT-S [29] 43 268 48.4 70.5 52.6 43.9 67.6 47.5
+Ortho-S [25] 44 277 48.7 70.5 53.3 43.6 67.3 47.3
+SMT-S [34] 40 265 49.0 70.1 53.4 43.4 67.3 46.7
+CSWin-T [11] 42 279 49.0 70.7 53.7 43.6 67.9 46.6
+InternImage-T [56] 49 270 49.1 70.4 54.1 43.7 67.3 47.3
+RMT-S 46 262 50.7 71.9 55.6 44.9 69.1 48.4
+ConvNeXt-S [36] 70 348 47.9 70.0 52.7 42.9 66.9 46.2
+NAT-S [19] 70 330 48.4 69.8 53.2 43.2 66.9 46.4
+Swin-S [35] 69 359 48.5 70.2 53.5 43.3 67.3 46.6
+InternImage-S [56] 69 340 49.7 71.1 54.5 44.5 68.5 47.8
+SMT-B [34] 52 328 49.8 71.0 54.4 44.0 68.0 47.3
+CSWin-S [11] 54 342 50.0 71.3 54.7 44.5 68.4 47.7
+RMT-B 73 373 52.2 72.9 57.0 46.1 70.4 49.9
+Table 3: Comparison to other backbones using Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
+Backbone Params (M) FLOPs (G) Cascade Mask R-CNN 3×3$\times$3 ×+MS
+$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
+Swin-T [35] 86 745 50.5 69.3 54.9 43.7 66.6 47.1
+NAT-T [19] 85 737 51.4 70.0 55.9 44.5 67.6 47.9
+GC-ViT-T [20] 85 770 51.6 70.4 56.1 44.6 67.8 48.3
+SMT-S [34] 78 744 51.9 70.5 56.3 44.7 67.8 48.6
+UniFormer-S [30] 79 747 52.1 71.1 56.6 45.2 68.3 48.9
+Ortho-S [25] 81 755 52.3 71.3 56.8 45.3 68.6 49.2
+HorNet-T [43] 80 728 52.4 71.6 56.8 45.6 69.1 49.6
+CSWin-T [11] 80 757 52.5 71.5 57.1 45.3 68.8 48.9
+RMT-S 83 741 53.2 72.0 57.8 46.1 69.8 49.8
+Swin-S [35] 107 838 51.9 70.7 56.3 45.0 68.2 48.8
+NAT-S [19] 108 809 51.9 70.4 56.2 44.9 68.2 48.6
+GC-ViT-S [20] 108 866 52.4 71.0 57.1 45.4 68.5 49.3
+DAT-S [58] 107 857 52.7 71.7 57.2 45.5 69.1 49.3
+HorNet-S [43] 108 827 53.3 72.3 57.8 46.3 69.9 50.4
+CSWin-S [11] 92 820 53.7 72.2 58.4 46.4 69.6 50.6
+UniFormer-B [30] 107 878 53.8 72.8 58.5 46.4 69.9 50.4
+RMT-B 111 852 54.5 72.8 59.0 47.2 70.5 51.4
+Table 4: Comparison to other backbones using Cascade Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
 Settings.
-We adopt MMDetection [3] to implement RetinaNet [29] and Mask-RCNN [20].
-We use the commonly used ”1×1$\times$1 ×” (12 training epochs) setting for the two strategies and ”$3\times+\mathrm{MS}$” for Mask-RCNN.
+We adopt MMDetection [4] to implement RetinaNet [32], Mask-RCNN [22] and Cascade Mask R-CNN [2].
+We use the commonly used “1×1$\times$1 ×” (12 training epochs) setting for the RetinaNet and Mask R-CNN.
+Besides, we use “$3\times+\mathrm{MS}$” for Mask R-CNN and Cascade Mask R-CNN.
 Following [35], during training, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels.
-We adopt the AdamW optimizer with a learning rate of 0.0001, weight decay of 0.05, and batch size of 16 to optimize the model.
-For ”1×1$\times$1 ×” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 8 and 11.
-While for ”$3\times+\mathrm{MS}$” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 27 and 33.
+We adopt the AdamW optimizer with a learning rate of 0.0001 and batch size of 16 to optimize the model.
+For the “1×1$\times$1 ×” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 8 and 11.
+While for the “$3\times+\mathrm{MS}$” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 27 and 33.
 Results.
 Tab.
+2, Tab.
 3 and Tab.
-3 shows the results with RetinaNet and Mask R-CNN.
+4 show the results with different detection frameworks.
 The results demonstrate that our RMT performs best in all comparisons.
-For the RetinaNet framework, our RMT-T outperforms FAT-B2 by +1.1 AP, while S/B/L also perform better than other methods.
-As for the Mask R-CNN with ”1×1$\times$1 ×” schedule, RMT-L outperforms the recent InternImage-B by +1.8 box AP and +1.9 mask AP.
-For ”$3\times+\mathrm{MS}$” schedule, RMT-S outperforms InternImage-T for +1.6 box AP and +1.2 mask AP.
+For the RetinaNet framework, our RMT-T outperforms MPViT-XS by +1.3 AP, while S/B/L also perform better than other methods.
+As for the Mask R-CNN with “1×1$\times$1 ×” schedule, RMT-L outperforms the recent InternImage-B by +2.8 box AP and +1.9 mask AP.
+For “$3\times+\mathrm{MS}$” schedule, RMT-S outperforms InternImage-T for +1.6 box AP and +1.2 mask AP.
+Besides, regarding the Cascade Mask R-CNN, our RMT still performs much better than other backbones.
 All the above results tell that RMT outperforms its counterparts by evident margins.
 4.3 Semantic Segmentation
-Backbone Semantic FPN
-Params(M) FLOPs(G) mIoU(%)
-DAT-T [51] 32 198 42.6
-CrossFormer-S [48] 34 221 46.0
-UniFormer-S [27] 25 247 46.6
-CSWin-T [9] 26 202 48.2
-Shuted-S [39] 26 183 48.2
-RMT-S 30 180 49.4
-DAT-S 53 320 46.1
-UniFormer-B [27] 54 350 47.7
-CrossFormer-B [48] 56 331 47.7
-CSWin-S [9] 39 271 49.2
-RMT-B 57 294 50.4
-DAT-B 92 481 47.0
-CrossFormer-L 95 497 48.7
-CSWin-B [9] 81 464 49.9
-RMT-L 98 482 51.4
-Table 4: Comparison with the state-of-the-art on ADE20K. The framework is SemanticFPN.
-Backbone UperNet
-Params(M) FLOPs(G) mIoU(%)
-DAT-T [51] 60 957 45.5
-NAT-T [18] 58 934 47.1
-InternImage-T [49] 59 944 47.9
-HorNet-T [38] 55 924 49.2
-SMT-S [31] 50 935 49.2
-RMT-S 56 937 49.8
-DAT-S [51] 81 1079 48.3
-SMT-B [31] 62 1004 49.6
-HorNet-S [38] 85 1027 50.0
-InterImage-S [49] 80 1017 50.2
-CSWin-S [9] 65 1027 50.4
-RMT-B 83 1051 51.5
-Table 5: Comparison with the state-of-the-art on ADE20K. The framework is UperNet.
+Backbone Method Params(M) FLOPs(G) mIoU(%)
+ResNet18 [21] FPN 15.5 32.2 32.9
+PVTv2-B1 [54] FPN 17.8 34.2 42.5
+VAN-B1 [17] FPN 18.1 34.9 42.9
+EdgeViT-S [38] FPN 16.9 32.1 45.9
+RMT-T FPN 17.0 33.7 46.4
+DAT-T [58] FPN 32 198 42.6
+RegionViT-S+ [3] FPN 35 236 45.3
+CrossFormer-S [55] FPN 34 221 46.0
+UniFormer-S [30] FPN 25 247 46.6
+Shuted-S [44] FPN 26 183 48.2
+RMT-S FPN 30 180 49.4
+DAT-S [58] FPN 53 320 46.1
+RegionViT-B+ [3] FPN 77 459 47.5
+UniFormer-B [30] FPN 54 350 47.7
+CrossFormer-B [55] FPN 56 331 47.7
+CSWin-S [11] FPN 39 271 49.2
+RMT-B FPN 57 294 50.4
+DAT-B [58] FPN 92 481 47.0
+CrossFormer-L [55] FPN 95 497 48.7
+CSWin-B [11] FPN 81 464 49.9
+RMT-L FPN 98 482 51.4
+DAT-T [58] UperNet 60 957 45.5
+NAT-T [19] UperNet 58 934 47.1
+InternImage-T [56] UperNet 59 944 47.9
+MPViT-S [29] UperNet 52 943 48.3
+SMT-S [34] UperNet 50 935 49.2
+RMT-S UperNet 56 937 49.8
+DAT-S [58] UperNet 81 1079 48.3
+SMT-B [34] UperNet 62 1004 49.6
+HorNet-S [43] UperNet 85 1027 50.0
+InterImage-S [56] UperNet 80 1017 50.2
+MPViT-B [29] UperNet 105 1186 50.3
+CSWin-S [11] UperNet 65 1027 50.4
+RMT-B UperNet 83 1051 52.0
+Swin-B [35] UperNet 121 1188 48.1
+GC ViT-B [20] UperNet 125 1348 49.2
+DAT-B [58] UperNet 121 1212 49.4
+InternImage-B [56] UperNet 128 1185 50.8
+CSWin-B [11] UperNet 109 1222 51.1
+RMT-L UperNet 125 1241 52.8
+Table 5: Comparison with the state-of-the-art on ADE20K.
 Settings.
@@
 The learning rate is set to $6\times 10^{-5}$ with 1500 iterations warmup.
+Model Params(M) FLOPs(G) Top1-acc(%) $AP^{b}$ $AP^{m}$ mIoU(%)
+DeiT-S [49] 22 4.6 79.8
+RMT-DeiT-S 22 4.6 81.7(+1.9)
+Swin-T [35] 29 4.5 81.3 43.7 39.8 44.5
+RMT-Swin-T 29 4.7 83.6(+2.3) 47.8(+4.1) 43.1(+3.3) 49.1(+4.6)
+Swin-S [35] 50 8.8 83.0 45.7 41.1 47.6
+RMT-Swin-S 50 9.1 84.5(+1.5) 49.5(+3.8) 44.2(+3.1) 51.0 (+3.4)
+RMT-T 14.3 2.5 82.4 47.1 42.6 46.4
+$\xrightarrow{}$ttention 14.3 2.5 81.6(-0.8) 44.6(-2.5) 40.7(-1.9) 43.9(-2.5)
+$\xrightarrow{}$ate 15.6 2.7 Nan
+w/o LCE 14.2 2.4 82.1 46.7 42.3 46.0
+w/o CPE 14.3 2.5 82.2 47.0 42.4 46.4
+w/o Stem 14.3 2.2 82.2 46.8 42.3 46.2
+Table 6: Ablation study. We make a strict comparison among RMT, DeiT, and Swin-Transformer.
+3rd stage FLOPs(G) Top1(%) FLOPs(G) mIoU(%)
+MaSA-d 4.5 84.1 180 49.4
+MaSA 4.8 84.1 246 49.7
+Table 7: Comparison between decomposed MaSA (MaSA-d) and original MaSA.
+[…]
+Parallel 27 10.9 262
+Chunklen_4 27 4.5 192
+Chunklen_49 27 4.7 446 82.1
+Recurrent 27 4.5 61
+MaSA 27 4.5 876 84.1
+Table 8: Comparison between MaSA and retention in RMT-S’s architecture.
+[…]
+BiFormer-T [75] 13 2.2 1602 81.4
+CMT-XS [16] 15 1.5 1476 81.8
+SMT-T [34] 12 2.4 636 82.2
+RMT-T 14 2.5 1650 82.4
+CMT-S [16] 25 4.0 848 83.5
+MaxViT-T [51] 31 5.6 826 83.6
+SMT-S [34] 20 4.8 356 83.7
+BiFormer-S [75] 26 4.5 766 83.8
+RMT-Swin-T 29 4.7 1192 83.6
+RMT-S 27 4.5 876 84.1
+SMT-B [34] 32 7.7 237 84.3
+BiFormer-B [75] 57 9.8 498 84.3
+CMT-B [16] 46 9.3 447 84.5
+MaxViT-S [51] 69 11.7 546 84.5
+RMT-Swin-S 50 9.1 722 84.5
+RMT-B 54 9.7 457 85.0
+SMT-L [34] 80 17.7 158 84.6
+MaxViT-B [51] 120 23.4 306 84.9
+RMT-L 95 18.2 326 85.5
+Table 9: Comparison of inference speed among SOTA models.
 Results.
 The results of semantic segmentation can be found in Tab.
-4 and Tab.
 5.
-All the FLOPs are measured with the resolution of $512\times 2048$.
+All the FLOPs are measured with the resolution of $512\times 2048$, except the group of RMT-T, which are measured with the resolution of $512\times 512$.
 All our models achieve the best performance in all comparisons.
 Specifically, our RMT-S exceeds Shunted-S for +1.2 mIoU with Semantic FPN.
-Moreover, our RMT-B outperforms the recent InternImage-S for +1.3 mIoU.
+Moreover, our RMT-B outperforms the recent InternImage-S for +1.8 mIoU.
 All the above results demonstrate our model’s superiority in dense prediction.
 4.4 Ablation Study
-Model Params(M) FLOPs(G) Top1-acc(%)
-RMT-T 14 2.5 82.4
-w/o $\gamma$ 14 2.5 81.6
-Table 6: Ablation study
-$\gamma$ decay.
-We verify the impact of explicit decay on the model, as shown in the Tab. 6. Explicit decay improves the performance of the model.
+Strict comparison with previous works.
+In order to make a strict comparison with previous methods, we align RMT’s hyperparameters (such as whether to use hierarchical structure, the number of channels in the four stages of the hierarchical model, whether to use positional encoding and convolution stem, etc.) of the overall architecture with DeiT [49] and Swin [35], and only replace the Self-Attention/Window Self-Attention with our MaSA.
+The comparison results are shown in Tab.
+6, where RMT significantly outperforms DeiT-S, Swin-T, and Swin-S.
+MaSA.
+We verify the impact of Manhattan Self-Attention on the model, as shown in the Tab.
+6.
+MaSA improves the model’s performance in image classification and downstream tasks by a large margin.
+Specifically, the classification accuracy of MaSA is 0.8% higher than that of vanilla attention.
+Softmax.
+In RetNet, Softmax is replaced with a non-linear gating function to accommodate its various computational forms [46].
+We replace the Softmax in MaSA with this gating function.
+However, the model utilizing the gating function cannot undergo stable training.
+It is worth noting that this does not mean the gating function is inferior to Softmax.
+The gating function may just not be compatible with our decomposed form or spatial decay.
+LCE.
+Local Context Enhancement also plays a role in the excellent performance of our model. LCE improves the classification accuracy of RMT by 0.3% and enhances the model’s performance in downstream tasks.
+CPE.
+Just like previous methods, CPE provides our model with flexible position encoding and more positional information, contributing to the improvement in the model’s performance in image classification and downstream tasks.
+Convolutional Stem.
+The initial convolutional stem of the model provides better local information, thereby further enhancing the model’s performance on various tasks.
+Decomposed MaSA.
+In RMT-S, we substitute the decomposed MaSA (MaSA-d) in the third stage with the original MaSA to validate the effectiveness of our decomposition method, as illustrated in Tab.
+7.
+In terms of image classification, MaSA-d and MaSA achieve comparable accuracy.
+However, for semantic segmentation, employing MaSA-d significantly reduces computational burden while yielding similar result.
+MaSA v.s. Retention.
+As shown in Tab.
+8, we replace MaSA with the original retention in the architecture of RMT-S.
+We partition the tokens into chunks using the method employed in Swin-Transformer [35] for chunk-wise retention.
+Due to the limitation of retention in modeling one-dimensional causal data, the performance of the vision backbone based on it falls behind RMT.
+Moreover, the chunk-wise and recurrent forms of retention disrupt the parallelism of the vision backbone, resulting in lower inference speed.
+Inference Speed.
+We compare the RMT’s inference speed with the recent best performing vision backbones in Tab. 9. Our RMT demonstrates the optimal trade-off between speed and accuracy.
 5 Conclusion
-We propose RMT, a vision backbone that integrates retentive network and Vision Transformer.
-RMT introduces explicit decay related to distance, which brings spatial prior knowledge to visual models.
-The new mechanism is called Retentive Self-Attention (ReSA).
-To reduce the complexity of the model, RMT also employs a method that decomposes the ReSA into two axes.
-Extensive experiments validate the effectiveness of RMT, especially in downstream tasks such as object detection, where RMT demonstrates significant advantages.
-Our work is still in progress.
+In this work, we propose RMT, a vision backbone with explicit spatial prior.
+RMT extends the temporal decay used for causal modeling in NLP to the spatial level and introduces a spatial decay matrix based on the Manhattan distance.
+The matrix incorporates explicit spatial prior into the Self-Attention.
+Additionally, RMT utilizes a Self-Attention decomposition form that can sparsely model global information without disrupting the spatial decay matrix.
+The combination of spatial decay matrix and attention decomposition form enables RMT to possess explicit spatial prior and linear complexity.
+Extensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate the superiority of RMT.
+Appendix A Architecture Details
+Model Blocks Channels Heads Ratios Params(M) FLOPs(G)
+RMT-T [2, 2, 8, 2] [64, 128, 256, 512] [4, 4, 8, 16] [3, 3, 3, 3] 14 2.5
+RMT-S [3, 4, 18, 4] [64, 128, 256, 512] [4, 4, 8, 16] [4, 4, 3, 3] 27 4.5
+RMT-B [4, 8, 25, 8] [80, 160, 320, 512] [5, 5, 10, 16] [4, 4, 3, 3] 54 9.7
+RMT-L [4, 8, 25, 8] [112, 224, 448, 640] [7, 7, 14, 20] [4, 4, 3, 3] 95 18.2
+RMT-DeiT-S [12] [384] [6] [4] 22 4.6
+RMT-Swin-T [2, 2, 6, 2] [96, 192, 384, 768] [3, 6, 12, 24] [4, 4, 4, 4] 29 4.7
+RMT-Swin-S [2, 2, 18, 2] [96, 192, 384, 768] [3, 6, 12, 24] [4, 4, 4, 4] 50 9.1
+Table 10: Detailed Architectures of our models.
+Our architectures are illustrated in the Tab.
+10.
+For convolution stem, we apply five $3\times 3$ convolutions to embed the image into $56\times 56$ tokens.
+GELU and batch normalization are used after each convolution except the last one, which is only followed by batch normalization.
+$3\times 3$ convolutions with stride 2 are used between stages to reduce the feature map’s resolution.
+$3\times 3$ depth-wise convolutions are adopted in CPE.
+Moreover, $5\times 5$ depth-wise convolutions are adopted in LCE.
+RMT-DeiT-S, RMT-Swin-T, and RMT-Swin-S are models that we used in our ablation experiments.
+Their structures closely align with the structure of DeiT [49] and Swin-Transformer [35] without using techniques like convolution stem, CPE, and others.
+Appendix B Experimental Settings
+ImageNet Image Classification.
+We adopt the same training strategy with DeiT [49] with the only supervision is the classification loss.
+In particular, our models are trained from scratch for 300 epochs.
+We use the AdamW optimizer with a cosine decay learning rate scheduler and 5 epochs of linear warm-up.
+The initial learning rate, weight decay, and batch size are set to 0.001, 0.05, and 1024, respectively.
+Our augmentation settings are RandAugment [8] (randm9-mstd0.5-inc1), Mixup [70] (prob=0.8), CutMix [69] (probe=1.0), Random Erasing [73] (prob=0.25) and Exponential Moving Average (EMA) [40].
+The maximum rates of increasing stochastic depth [24] are set to 0.1/0.15/0.4/0.5 for RMT-T/S/B/L, respectively.
+For a more comprehensive comparison, we train two versions of the model.
+The first version uses only classification loss as the supervision, while the second version, in addition to the classification loss, incorporates token labeling introduced by [27] for additional supervision.
+Models using token labeling are marked with“*”.
+COCO Object Detection and Instance Segmentation.
+We apply RetinaNet [32], Mask-RCNN [22] and Cascaded Mask-CNN [2] as the detection frameworks to conduct experiments.
+We implement them based on the MMDetection [4].
+All models are trained under two common settings:“1×1$\times$1 ×” (12 epochs for training) and“3×3$\times$3 ×+MS” (36 epochs with multi-scale augmentation for training).
+For the “1×1$\times$1 ×” setting, images are resized to the shorter side of 800 pixels.
+For the “3×3$\times$3 ×+MS”, we use the multi-scale training strategy and randomly resize the shorter side between 480 to 800 pixels.
+We apply AdamW optimizer with the initial learning rate of 1e-4.
+For RetinaNet, we use the weight decay of 1e-4 for RetinaNet while we set it to 5e-2 for Mask-RCNN and Cascaded Mask-RCNN.
+For all settings, we use the batch size of 16, which follows the previous works [35, 63, 64]
+ADE20K Semantic Segmentation.
+Based on MMSegmentation [7], we implement UperNet [59] and SemanticFPN [28] to validate our models.
+For UperNet, we follow the previous setting of Swin-Transformer [35] and train the model for 160k iterations with the input size of $512\times 512$.
+For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.
+Appendix C Efficiency Comparison
+We compare the inference speed of RMT with other backbones, as shown in Tab. 11. Our models achieve the best trade-off between speed and accuracy among many competitors.
+Model Params (M) FLOPs (G) Troughput (imgs/s) Top1 (%)
+MPViT-XS [29] 11 2.9 1496 80.9
+Swin-T [35] 29 4.5 1704 81.3
+BiFormer-T [75] 13 2.2 1602 81.4
+GC-ViT-XT [20] 20 2.6 1308 82.0
+SMT-T [34] 12 2.4 636 82.2
+RMT-T 14 2.5 1650 82.4
+Focal-T [63] 29 4.9 582 82.2
+CSWin-T [11] 22 4.3 1561 82.7
+Eff-B4 [47] 19 4.2 627 82.9
+MPViT-S [29] 23 4.7 986 83.0
+Swin-S [35] 50 8.8 1006 83.0
+SGFormer-S [15] 23 4.8 952 83.2
+iFormer-S [45] 20 4.8 1051 83.4
+CMT-S [16] 25 4.0 848 83.5
+RMT-Swin-T 29 4.7 1192 83.6
+CSwin-S [11] 35 6.9 972 83.6
+MaxViT-T [51] 31 5.6 826 83.6
+SMT-S [34] 20 4.8 356 83.7
+BiFormer-S [75] 26 4.5 766 83.8
+RMT-S 27 4.5 876 84.1
+Model Params (M) FLOPs (G) Troughput (imgs/s) Top1 (%)
+Focal-S [63] 51 9.1 351 83.5
+Eff-B5 [47] 30 9.9 302 83.6
+SGFormer-M [15] 39 7.5 598 84.1
+SMT-B [34] 32 7.7 237 84.3
+BiFormer-B [75] 57 9.8 498 84.3
+RMT-Swin-S 50 9.1 722 84.5
+MaxViT-S [51] 69 11.7 546 84.5
+CMT-B [16] 46 9.3 447 84.5
+iFormer-B [45] 48 9.4 688 84.6
+RMT-B 54 9.7 457 85.0
+Swin-B [35] 88 15.5 756 83.5
+Eff-B6 [47] 43 19.0 172 84.0
+Focal-B [63] 90 16.4 256 84.0
+CSWin-B [11] 78 15.0 660 84.2
+MPViT-B [29] 75 16.4 498 84.3
+SMT-L [34] 80 17.7 158 84.6
+SGFormer-B [15] 78 15.6 388 84.7
+iFormer-L [45] 87 14.0 410 84.8
+MaxViT-B [51] 120 23.4 306 84.9
+RMT-L 95 18.2 326 85.5
+Table 11: Comparison of inference speed.
+Appendix D Details of Explicit Decay
+We use different $\gamma$ for each head of the multi-head ReSA to control the receptive field of each head, enabling the ReSA to perceive multi-scale information.
+We keep all the $\gamma$ of ReSA’s heads within a certain range.
+Assuming the given receptive field control interval of a specific ReSA module is $[a,b]$, where both $a$ and $b$ are positive real numbers.
+And the total number of the ReSA module’s heads is $N$.
+The $\gamma$ for its $i$th head can be written as Eq.
+8:
+$\gamma_{i}=1-2^{-a-\frac{(b-a)i}{N}}$ (8)
+For different stages of different backbones, we use different values of $a$ and $b$, with the details shown in Tab. 12.
+Model $a$ $b$
+RMT-T [2, 2, 2, 2] [6, 6, 8, 8]
+RMT-S [2, 2, 2, 2] [6, 6, 8, 8]
+RMT-B [2, 2, 2, 2] [7, 7, 8, 8]
+RMT-L [2, 2, 2, 2] [8, 8, 8, 8]
+RMT-DeiT-S [2] [8]
+RMT-Swin-T [2, 2, 2, 2] [8, 8, 8, 8]
+RMT-Swin-S [2, 2, 2, 2] [8, 8, 8, 8]
+Table 12: Details about the $\gamma$ decay.